# References.bib

### v1

@@ 6.1 Corpus on Blockchain
Large language models require massive amounts of data to operate, with modern corpus sizes typically exceeding dozens of terabytes in volume [zhao_survey_2023]. This inherent quality of LLMs is situated opposite to problems that blockchain systems typically prefer to solve. Reconciling the limitations of blockchain’s throughput and capacity for data handling with LLMs ballooning size is perhaps the most pressing issue in the BC4LLMs space.
Multiple attempts in which researchers have proposed zero-knowledge proofs for scalability have been discussed, such as in Wellington [wellington_basedai_2024] and Singh [singh_enhancing_2024]. However, employing zero-knowledge technology solely for scalability, and not privacy, is a doubtful long-term prospect due to the demands of generating zero-knowledge proofs, even with incredibly small circuits. Largely, this is due to the long-standing problem of the MSM within the context of generating ZKPs [xavier_pipemsm_2022]. Moreover, current WebGPU and/or WASM implementations could likely not support the throughput that client-based LLMs would require. For these reasons, it is unlikely that zero knowledge could be employed as a definitive solution for scalability issues within BC4LLMs without extensive developments in zk-SNARK generation research.
The disconnect between large language models’ growing size and the aversion of blockchain’s to store data readily on chain is a major research challenge.
@@ 6.2 Reliance on Oracles
The security guarantees of blockchain technology, while robust, often leave little room for interoperability with external systems [deshpande_distributed_nodate]. That is, the blockchain can most easily interact with information on the chain, leaving little room to consider issues such as fact-checking or moral alignment. Oftentimes, to develop mechanisms that seek to provide assistance with LLM toxicity or factuality, oracles are used to bridge this gap [fan_blockchain_2024][fraga-lamas_fake_2020]. Serving as mediators between chains and online sources, oracles are trusted parties that deliver information through a variety of protocols and frameworks. However, introducing a trusted party into an otherwise trustless system has been a long-standing weak point in this solution [mbula_mboma_assessing_2023]. Exploring non-oracle-based options for ground truth solutions, or toxicity checks, would greatly enhance the security guarantees of blockchain within LLMs.
@@ 6.3 Energy Consumption
The crux of many issues concerning large language models stems from their need to consume and process inordinate amounts of data [zhao_survey_2023]. This requirement, in turn, creates a corresponding need for large language models to consume equally as colossal quantities of energy during their training stages, as well was during run time [luo_bc4llm_2023]. On the other side of the coin, we have energy issues within blockchain as well. Oftentimes, blockchain systems struggle to limit energy costs as the demands of both consensus mechanisms and the validation of proposed transactions incur huge computational costs [mbula_mboma_assessing_2023]. Due to their tendency’s to consume large amounts of compute resources, we see an unfortunate dissonance between implementing BC4LLM at scale without large efforts to cut back on energy costs, possibly through a shift away from both transformer architectures and proof of work type consensus mechanisms [nguyen_proof-of-stake_2019].
@@ 7.1 Blockchain Federated Unlearning
Privacy regulations are paramount in the online realm, especially concerning the ”right to be forgotten” and user data privacy which are critical considerations when working with LLMs and blockchain. Federated blockchain unlearning offers LLMs the ability to erase learned data. Within our research, we identified four recent papers that have implemented blockchain federated unlearning frameworks. As noted previously in Section 4.1.1, Zuo et al. [zuo_federatedtrustchain_2024] developed a federated TrustChain framework for blockchain-enhanced LLM training and unlearning, focusing on the impact of Low-Rank Adaptation (LoRA) hyperparameters on unlearning performances and integrating Hyperledger Fabric to ensure the security, transparency, and verifiability of the unlearning process. In another study, Zuo et al. [zuo_federated_2024] presented a trustworthy approach towards federated learning with blockchain-enhanced machine unlearning. This implementation differs from Trustchain, where Zuo et al. [zuo_federated_2024] used a machine unlearning mechanism that utilized two types of clients for training and unlearning, smart contracts for process automation, and a blockchain network for secure, immutable record-keeping. Beyond the above works [zuo_federated_2024], Liu et al. [liu_decentralized_2024] introduced Blockchain Federated Unlearning (BlockFUL) as a versatile framework that redesigns the blockchain structure using a Chameleon Hash (CH) technology to simplify model updates and reduce the computational and consensus costs associated with unlearning tasks. Additionally, BlockFUL ensures the integrity and traceability of model updates, including privacy-preserving results from these blockchain-based unlearning operations [liu_decentralized_2024]. Lin et al. [lin_blockchain-enabled_2024] propose a framework with a proof of federated unlearning protocol that also utilizes the Chameleon hash function to verify data removal and eliminate the data contributions stored in other clients’ models. Both, Liu et al. and Lin et al. use CH functions in their blockchain-enabled federated unlearning processes. The applications of key blockchain components, such as on-chain smart contracts and hash mappings for verifying data removal, may enable LLMs to forget personal data effectively. Blockchain for unlearning is an emerging area of research with significant potential for further innovation.
7.2 Blockchain Enhanced Retrieval Augmented Generation
Considering the novelty of RAG, there exists a plethora of research that focuses on how RAG may be attacked, consequently weakening the integrity of large language models. [xue_badrag_2024][anderson_is_2024][deng_pandora_2024][zou_poisonedrag_2024][hu_prompt_2024][zeng_good_2024][cheng_trojanrag_2024] However, there exists a vacuum yet to be filled with potential work concerning how these attacks may be mitigated, especially where defenses can make use of blockchain technology. In particular, calls for the inspection of blockchain’s potential role in RAG deployment [balakrishnan_enhancing_2024] have been made, and preliminary work done in exploring this intersection where it concerns blockchain’s potential usefulness in user experience [zhang_interactive_2024] and performance assessment [park_blockchain-based_2021] has been conducted. However, efforts concentrated on the enhancement of security and safety within RAG systems are absent from the current literature. A concerted effort towards bettering BC4LLMs where it concerns RAG would be both mutually and independently beneficial for blockchain technology and large language models.
7.3 Blockchain for Privacy Guarantees in Large Language Models
@@
The clear connection between federated learning, blockchain, and large language models allows for the field of differential privacy to enter BC4LLMs’ sphere of relevance. Major contributions concerning the impact of differential privacy on related areas such as deep learning have already been made [abadi_deep-learning-dp], but issues such as privacy budget exhaustion still loom large in the space [bkakria_budget-dp]. Moreover, despite conclusions that blockchain can help with privacy budget exhaustion [han_blockchain-based_2020, zhao_blockchain-based_2021], few efforts have been conducted in exploring these solutions. Indeed, there is a need for more relevant research in order to realize the full measure of blockchain’s impact on this area.
7.4 Blockchain for Transparency and Data Provenance in Large Language Models
@@
Several recent papers have urged for increased data accountability measures to be placed on organizations developing large language models, especially where it concerns issues of data acquisition [bender_dangers_2021][he_survey_2024]. Additionally, worth noting are direct calls for the introduction of blockchain technology to help solve the issue of data provenance [werder_establishing_2022] in large language models. Largely, while this has been answered with responses in the realms of auditability [li_blockchain-based_2023], straightforward data tracking solutions have remained absent from the literature, despite relatively simple conceptual formulations [dinh_ai_2018]. Towards this goal of achieving improved data provenance within large language models corpus’, RAG databases, and even in-context learning repositories, there is a need for more explorations into this natural application of distributed ledger technology to problems of explainable AI concerning large language models.
7.5 Blockchain for Non-toxic Large Language Models
@@
Encompassing vital attributes such as ethics, legality, and non-violence, developing non-toxic large language models has been and will continue to be a major focus of the field for the foreseeable future [liu_trustworthy_2024][han_towards_2024][shen_large_2023]. There is no doubt that automated filtering of generated toxic content is one of the most pressing challenges concerning the safety of large language models [gehman_realtoxicityprompts_2020][anwar_foundational_2024]. This is because in essence, filtering inferences negatively impacts the quality of LLM responses, whereas manual human annotation is a costly and complex process [anwar_foundational_2024]. Therefore, the applications of blockchain technology in this regard, while currently limited, are compelling. Considering one of the most groundbreaking achievements in ML within the past several years, federated learning has allowed for massive strides to be made within the spaces of securing training sets, user privacy, and even misinformation defense. A similar approach, aimed at the problem of toxicity, could be a hugely beneficial endeavor to the field. Moreover, imagining such a model is not difficult. Developing consensus around what is considered correct in a model and using that to propagate gradients and parameters is not dissimilar to the decisions that must be made about what is or is not toxic given the state of certain corpora. Given a concentrated research program, automated non-toxicity could very well have excellent solutions founded within the blockchain space.
@@ 8 Conclusion
In this survey, we presented how large language models have been applied to various facets of everyday life, revealing significant system vulnerabilities such as data poisoning, hallucinations, jailbreaking, and privacy attacks. Despite widespread recognition of these issues, effective mitigation strategies remain limited. Traditional machine learning solutions, such as differential privacy and federated learning have been applied, but fall short in providing comprehensive protection against the unique threats faced by LLMs. The adoption of blockchain technology presents a promising avenue for enhancing the security and safety of LLMs. Blockchain systems offer robust mechanisms for ensuring data integrity, provenance and encrypted frameworks which can be leveraged to bolster LLM defenses. By incorporating blockchain-based defenses, it is viable to achieve stronger privacy preservation, and reliable data, and offer LLMs to be more resilient against adversarial threats.
Furthermore, it is critical to establish clear definitions of security and safety in the context of LLM technologies. We conclude that security for LLMs pertains to the ability to tolerate applicable adversarial attacks while maintaining system integrity to provide consistent and accurate responses. Whereas safety for LLMs is the model’s capacity to interact with users in a trustworthy manner, contingent upon adhering to ethical concerns, law-abiding, non-violent, fair, passively privacy-preserving, and informing. Additionally, differentiating between active and passive privacy measures will aid in developing more targeted and effective privacy-preserving strategies. These distinctions and definitions provide a foundational framework for future research in the BC4LLM space. From analyzing the integration of blockchain technology and LLMs, we provided a taxonomy in Figure 3 where previous research done in the field of BC4LLMs can apply to security and safety problems that LLMs face. We recognize various gaps in the BC4LLM space that need to be looked into for further consideration. In refining our understanding of relevant concepts, we see that the intersection of blockchain technology and LLMs holds significant potential for addressing the current shortcomings in LLM security and safety. With our contributions, we can use BC4LLM to provide more secure, reliable, and safe AI systems.

### v2

@@ 6.1 Corpus on Blockchain
LLMs’ training heavily relies on substantial data, with modern corpora typically exceeding dozens of terabytes in volume [zhao_survey_2023]. This characteristic is inherently misaligned with the constraints that blockchain systems are generally designed to address. Reconciling blockchain’s limitations in throughput and data-handling capacity with the extensive data requirements of LLMs represents one of the most pressing challenges in BC4LLMs. Several approaches have explored the use of zero-knowledge proofs (ZKPs) to enhance scalability [wellington_basedai_2024, singh_enhancing_2024]. However, relying on zero-knowledge technology solely for scalability, and not privacy, poses challenges due to the computational cost of generating ZKPs, even with minimal circuits. A significant factor here is the ongoing issue of the Multi-Scalar Multiplication (MSM) in ZKP generation [xavier_pipemsm_2022]. Furthermore, current WebGPU and WASM implementations likely fall short of the throughput required by client-based LLMs. For these reasons, it is improbable that zero-knowledge could serve as a definitive solution to scalability in BC4LLMs without significant advancements in zk-SNARK generation research. Addressing the discrepancy between the growing size of LLMs and blockchain’s limited capacity for on-chain data storage remains a substantial research challenge.
@@ 6.2 Reliance on Oracles
The security guarantees of blockchain technology, while robust, often leave little room for interoperability with external systems [deshpande_distributed_nodate]. That is, the blockchain can most easily interact with information on the chain, leaving little room to consider issues such as fact-checking or moral alignment. Oftentimes, to develop mechanisms that seek to assist with LLM toxicity or factuality, oracles are used to bridge this gap [fraga-lamas_fake_2020, fan_blockchain_2024]. Serving as mediators between chains and online sources, oracles are trusted parties that deliver information through a variety of protocols and frameworks. However, introducing a trusted party into an otherwise trustless system has been a long-standing weak point in this solution [mbula_mboma_assessing_2023]. Exploring non-oracle-based options for ground truth solutions, or toxicity checks, would greatly enhance the security guarantees of blockchain within LLMs.
@@ 6.3 Energy Consumption
A significant portion of the challenges associated with LLMs arises from their need to consume and process vast amounts of data [zhao_survey_2023]. This requirement, in turn, necessitates extensive energy consumption during both training and runtime [luo_bc4llm_2023]. On the other hand, blockchain systems face their own energy challenges, as consensus mechanisms and transaction validation processes often incur substantial computational costs [mbula_mboma_assessing_2023]. The high computational demands of both LLMs and blockchain systems highlight a misalignment with the scalability of BC4LLM implementations without substantial efforts to reduce energy costs. This might require moving away from transformer-based architectures and energy-intensive consensus mechanisms, such as proof of work, toward more sustainable alternatives [nguyen_proof-of-stake_2019].
@@ 7.1 Blockchain Federated Unlearning
Privacy regulations are paramount in the online realm, especially concerning the “right to be forgotten” and user data privacy which are critical considerations when working with LLMs and blockchain. Federated blockchain unlearning offers LLMs the ability to erase learned data. Within our research, we identified four recent papers that have implemented blockchain-based federated unlearning frameworks. As noted previously in Section 4.1.1, Zuo et al. [zuo_federatedtrustchain_2024] develop a federated TrustChain framework for blockchain-enhanced LLM training and unlearning, focusing on the impact of Low-Rank Adaptation (LoRA) hyperparameters on unlearning performances and integrating Hyperledger Fabric to ensure the security, transparency, and verifiability of the unlearning process. In another study, Zuo et al. [zuo_federated_2024] presented a trustworthy approach towards federated learning with blockchain-enhanced machine unlearning. This implementation differs from Trustchain, where Zuo et al. [zuo_federated_2024] used a machine unlearning mechanism that utilized two types of clients for training and unlearning, smart contracts for process automation, and a blockchain network for secure, immutable record-keeping. Beyond the above works [zuo_federated_2024], Liu et al. [liu_decentralized_2024] introduced Blockchain Federated Unlearning (BlockFUL) as a versatile framework that redesigns the blockchain structure using a Chameleon Hash (CH) technology to simplify model updates and reduce the computational and consensus costs associated with unlearning tasks. Additionally, BlockFUL ensures the integrity and traceability of model updates, including privacy-preserving results from these blockchain-based unlearning operations [liu_decentralized_2024]. Lin et al. [lin_blockchain-enabled_2024] propose a framework with a proof of federated unlearning protocol that also utilizes the Chameleon hash function to verify data removal and eliminate the data contributions stored in other clients’ models. Both use CH functions in their blockchain-enabled federated unlearning processes. The applications of key blockchain components, such as on-chain smart contracts and hash mappings for verifying data removal, may enable LLMs to forget personal data effectively. Blockchain for unlearning is an emerging area of research with significant potential for further innovation.
7.2 Blockchain-enhanced RAG
Considering the popularity of Retrieval Augmented Generation (RAG), extensive studies have emerged to focus on potential vulnerabilities within RAG that could compromise the integrity of LLMs [xue_badrag_2024, anderson_is_2024, deng_pandora_2024, zou_poisonedrag_2024, hu_prompt_2024, zeng_good_2024, cheng_trojanrag_2024]. However, a significant gap remains in addressing strategies to mitigate these attacks, particularly where blockchain technology could offer defensive benefits. Recent studies have called for exploring blockchain’s role in RAG deployment [balakrishnan_enhancing_2024], and preliminary investigations have assessed blockchain’s potential to enhance user experience [zhang_interactive_2024] and performance evaluation [park_blockchain-based_2021]. Nonetheless, dedicated efforts to strengthen security and safety within RAG systems are largely absent in current literature. Advancing BC4LLMs specifically in the context of RAG security could yield considerable mutual benefits for both blockchain and LLM technologies.
7.3 Blockchain for Privacy Guarantees in LLMs
@@
The clear connection between federated learning, blockchain, and LLMs allows for the field of differential privacy to enter BC4LLMs’ sphere of relevance. Major contributions concerning the impact of differential privacy on related areas such as deep learning have already been made [abadi_deep-learning-dp], but issues such as privacy budget exhaustion still loom large in the space [bkakria_budget-dp]. Moreover, despite conclusions that blockchain can help with privacy budget exhaustion [han_blockchain-based_2020, zhao_blockchain-based_2021], few efforts have been conducted in exploring these solutions. Indeed, there is a need for more relevant research in order to realize the full measure of blockchain’s impact on this area.
7.4 Blockchain for Data Provenance and Transparency in LLMs
@@
Several recent papers have urged for increased data accountability measures to be placed on organizations developing LLMs, especially where it concerns issues of data acquisition [bender_dangers_2021, he_survey_2024]. Additionally, worth noting are direct calls for the introduction of blockchain technology to help solve the issue of data provenance [werder_establishing_2022] in LLMs. Largely, while this has been answered with responses in the realms of auditability [li_blockchain-based_2023], straightforward data tracking solutions have remained absent from the literature, despite relatively simple conceptual formulations [dinh_ai_2018]. Towards this goal of achieving improved data provenance within LLMs corpus’, RAG databases, and even in-context learning repositories, there is a need for more explorations into this natural application of distributed ledger technology to problems of explainable AI concerning LLMs.
7.5 Blockchain for Non-toxic LLMs
@@
Encompassing vital attributes such as ethics, legality, and non-violence, developing non-toxic LLMs has been and will continue to be a major focus of the field for the foreseeable future [shen_large_2023, liu_trustworthy_2024, han_towards_2024]. There is no doubt that automated filtering of generated toxic content is one of the most pressing challenges concerning the safety of LLMs [gehman_realtoxicityprompts_2020, anwar_foundational_2024]. This is because in essence, filtering inferences negatively impacts the quality of LLM responses, whereas manual human annotation is a costly and complex process [anwar_foundational_2024]. Therefore, the applications of blockchain technology in this regard, while currently limited, are compelling. Considering one of the most groundbreaking achievements in ML within the past several years, federated learning has allowed for massive strides to be made within the spaces of securing training sets, user privacy, and even misinformation defense. A similar approach, aimed at the problem of toxicity, could be a hugely beneficial endeavor to the field. Moreover, imagining such a model is not difficult. Developing consensus around what is considered correct in a model and using that to propagate gradients and parameters is not dissimilar to the decisions that must be made about what is or is not toxic given the state of certain corpora. Given a concentrated research program, automated non-toxicity could very well have excellent solutions found in the blockchain space.
@@ 8 Conclusion
In this survey, we first highlight significant systemic vulnerabilities in large language models (LLMs), including data poisoning, hallucinations, jailbreaking, and privacy attacks. Although these issues have been extensively studied in conventional machine learning models, with approaches like differential privacy and federated learning, comprehensive protection for LLMs remains an area for improvement. In contrast, blockchain technology offers a promising solution to enhance the security and safety of LLMs. Blockchain systems provide powerful mechanisms to ensure data integrity, provenance, and encrypted frameworks, which can be leveraged to strengthen LLM defenses. By integrating blockchain-based defenses, it is possible to achieve stronger privacy protection, reliable data, and improved resilience of LLMs against adversarial threats.
Besides, it is critical to establish clear definitions of security and safety in the context of LLMs. We conclude that security for LLMs pertains to the ability to tolerate applicable adversarial attacks while maintaining system integrity to provide consistent and accurate responses, whereas safety for LLMs is the model’s capacity to interact with users in a trustworthy manner, contingent upon adhering to ethical concerns, law-abiding, non-violent, fair, passively privacy-preserving, and informing. Additionally, differentiating between active and passive privacy measures will aid in developing more targeted and effective privacy-preserving strategies. These distinctions and definitions provide a foundational framework for future research in BC4LLM. From analyzing the integration of blockchain and LLMs, we propose a new taxonomy in Figure 3, where previous research done in the field of BC4LLMs can apply to security and safety problems that LLMs face. We recognize various gaps in BC4LLMs that need to be looked into for further consideration. In refining our understanding of relevant concepts, we see that the intersection of blockchain and LLMs holds significant potential for addressing the current shortcomings in LLM security and safety. Through our review, we aim to guide new researchers in understanding how blockchain technology can be utilized to enhance the security, reliability, and safety of LLMs.